Home‎ > ‎Crew 136‎ > ‎

227-CommandersReport

Commander Report
02/27/2014
Corentin Liber

Today we experienced bad weather as expected from the forecast and we had to change some of our plans.

We could luckily observe a little in the MUSK Observatory tonight between the clouds.  We were able
to record the spectrum of Jupiter, Sirius ans Arcturus.  Those spectrum are currently processed and
analyzed.  The Geologist also checked his samples under UV light during the night.

In the morning, we worked on individual chores and had some improvement in our streambed modeling. 
The conclusions will soon be able to be written on this project.

We had an ATV EVA planned in the afternoon but we had to abort because of the rain.  The EVA team
came back soon after leaving the Hab seeing the rainy weather coming fast toward the MDRS.

We expect some more rain in the evening but we still hope to be able to go out and record some more
spectrum this night.

--------------------------

CommanderCheckIn Report
02/27/2014
Thibaud Herriau

Crew Physical Health:
All are well.
Eva Depart Time:
1:10 pm
Eva Return Time:
1:45 pm
Narrative Of Field Mission Results:
We took the ATVs to go to the target area.
Unfortunately, we had to turn back after 10minutes due to the upcoming storm.
We all arrived home safely.

Hab Maint:
None
Rep Sched:
None
Tomorrow Plan:
EVA on foot to the North for the streambed project.
Wrapping up all projects and science reports.
Inventory:
None
Support Req:
None
Miscelleneous:
None